A baby weighing 2790 grams at birth is how many pounds/ounces?
Answer by checkley77(12844)

1 Pound=453.5923 Grams
1/453.5923=x/2,790
453.5923x=2,790
x=2,790/453.5923
x=6.15 Pounds or 6 pounds 2.4 Ounces.